Mathematics 134 – Intensive Calculus for Science 2 Quiz 3 – February 24, 2006

Your Name:

Do all work on this sheet (front and back if needed). You have 15 minutes. No calculators on this quiz.

A) (15) Integrate by parts, showing all work, choice of u and dv, and so forth:

∫ (x + 3)e^{2x} dx

(Note: you can check your work with the table.)

B) (10) Integrate using the table – say which entry you are using:

∫ \frac{dx}{x^{2} − 6x + 8}